An 8.00g sample occupies 22.4 L at 2.00 atm and 273 K. What is the molar mass of the gas?
1
Expert's answer
2017-04-14T06:50:19-0400
Solution.
PV = nRT
n = PV/RT
n = (2.00 atm x 22.4 L) / (0.08206 L atm / K mol x 273 K) = 2.00 mol
Molar mass = m/n = 8.00 g / 2.00 mol = 4 g/mol


Answer: M = 4 g/mol